write an equation for the line that passes through (-2,3) and is parallel to the line whos equation is 3x+2y=6
Answer by stanbon(75887)

write an equation for the line that
passes through (-2,3) and is parallel
to the line whos equation is 3x+2y = 6
--------
The given line has slope = -3/2
------------------
Equation Form: y = mx + b
You know y = 3 when x = -2 and m = -3/2; solve for "b":
-----
3 = (-3/2)(-2) + b
3 = 3 + b
b = 0
----
Equation:
y = (-3/2)x
================
Cheers,
Stan H.